Tom has 27 coins. They are all nickels and dimes. If he has a total of $1.95, how many nickels does he have?

Answer by fractalier(6550)

Let N be the number of nickels.
Let D be the number of dimes.
Thus N + D = 27
The value equation is 5N + 10D = 195
Now multiply the first equation by five and subtract it from the second...
5N + 10D = 195
-(5N + 5D = 135)
we get
5D = 60
D = 12
N = 15
To check, $1.20 + .75 = $1.95